Dale Decor sells home decor items through three distribution channels—retail stores, the Internet, and catalog sales. Each distribution channel is evaluated as an investment center. Selected results from the latest year are as follows: Retail Stores Internet Catalog Sales Sales revenue $10,000,000 $4,000,000 $3,200,000 Variable expenses 4,000,000 1,500,000 1,800,000 Direct fixed expenses 4,500,000 1,000,000 1,200,000 Average assets 8,000,000 4,000,000 2,000,000 Required rate of return 12% 12% 12% Collapse question part (a) Calculate the current residual income for each distribution channel. (If the residual income is a loss then enter with a negative sign preceding the number, e.g. -5,125 or parenthesis, e.g. (5,125).) Residual Income Retail $ Online $ Catalog $
Answer #1

Calculate residual income

Retail Online Catalog
Sales 10000000 4000000 3200000
Variable expense 4000000 1500000 1800000
Direct fixed expense 4500000 1000000 1200000
Operating income 1500000 1500000 200000
Average operating assets 8000000 4000000 2000000
Minimum return 960000 480000 240000
Residual income 1500000-960000 = 540000 1500000-480000 = 1020000 200000-240000 = -40000